OpenStudy (anonymous):

The volume of a cone is 565.2 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? (Use pie=3.14) 3 6 18 36

OpenStudy (anonymous):

do you know the formula for the volume of the cone?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

okay well google volume of a cone formula and find out :)

OpenStudy (anonymous):

The volume of a cone is 1/3(Area of Base)(height) = 1/3 π r2 h

OpenStudy (anonymous):

okay great so know plug in all the information you know into \[v = 1/3 \pi r^2 h\]

OpenStudy (anonymous):

hint v = 565.2 and h = 15

OpenStudy (anonymous):

now show me what you get

OpenStudy (mathstudent55):

Replace each variable with the shown value. Then solve for r^2.